# ConceptMaster: Multi-Concept Video Customization on Diffusion Transformer Models Without Test-Time Tuning

### v1

@@ 5.4 Multi-Concept Embeddings Injection Manner
In section 4.2, our ConceptMaster introduces a standalone MC-Injector to integrate the multi-concept multi-modal representation into the diffusion models. Some previous methods, represented as BLIP-Diffusion [33] and IP-Adapter [67], the former combines visual embeddings with textual caption embeddings as the whole condition representation, while the latter encodes the whole image as visual embeddings and aggregates into models by a decoupled cross-attention layer. However, in multi-concept scenarios, merging multi-modal features as the whole conditions can make it challenging to distinguish the semantic meanings among different identities. In addition, integrating all visual concepts on one image is not the optimal choice for multi-concept representation, especially when they contain similar visual appearances. We conduct the above two integration approaches of the multi-concept embeddings on our text-to-video generation models with the same training data for ConceptMaster. In Tab. 2 and Fig. 7, we can see that when customizing multiple concepts, both these two methods can hardly maintain the concept fidelity and deal with the identity decoupling problem. Additionally, when merging the textual and visual embeddings, the instruction following ability is also unsatisfactory. The generated video has a sense of splicing, and the dynamic degree is significantly reduced, as the original text cross-attention layer is influenced by the additional visual embeddings. Therefore, our designed ConceptMaster adopts the optimal solution to inject the decoupled multi-concept embeddings into the diffusion models.
@@ 5.5 Ablation Study
In section 4.2, our ConceptMaster proposes to firstly utilizes a Q-Former network to integrate the dense visual tokens extracted by CLIP image encoder into the comprehensive visual embeddings. While after simply replacing the Q-Former by an MLP layer, the generated videos cannot capture the appearances of given images, as in Fig. 8, and the quantitative metrics also largely drop in Tab. 3. In addition, in order to demonstrate the effectiveness of the designed DAM module, which is the intra-pair attention module based on paired visual embeddings and textual descriptions representation, we conduct several variants include: 1) Removing the DAM module, where only the extracted visual embeddings are further injected into the diffusion models, and the textual descriptions are unused. 2) Replacing the intra-pair attention by firstly concatenating the visual and textual embeddings along channel dimension (double the channel dimension), and integrating to the original channel dimension by an MLP layer. 3) Replacing the intra-pair cross-attention by self-attention, the features conduct self-attention are obtained by directly adding the visual and textual embeddings. In Tab. 3 and Fig. 8, we can see that the proposed DAM module is the optimal design. Cooperating the textual descriptions is significant, which not only enhances the uniqueness of each concept representation, but also assists the alignment of the multi-concept embeddings and the original diffusion model space. Furthermore, fusing visual and textual embeddings by the MLP layer is less comprehensive as it does not involve sufficient token-level interaction between them, lowering the concept representation and instruction-following in generated results. Additionally, the cross-attention operation is better than self-attention, which maintains consistent visual appearances in videos, while more artifacts (i.e., the cat’s head is split open) would be created by the self-attention operation.
@@ 6 Conclusion
In this paper, we introduce ConceptMaster, an innovative framework that effectively addresses the critical issues of identity decoupling while maintaining concept fidelity when customizing multiple identities in videos. ConceptMaster introduces a novel strategy for learning decoupled multi-concept embeddings and injecting them into diffusion models in a standalone manner. This strategy ensures the quality of customized videos with multiple identities, even for highly similar visual concepts. To further address the scarcity of high-quality multi-concept video-entity data, we have established a meticulous data construction pipeline. This pipeline enables the systematic collection of precise multi-concept video-entity data across diverse concepts. Additionally, we have designed a comprehensive benchmark to validate the effectiveness of our model from three critical dimensions across six different concept composition scenarios. Extensive experiments demonstrate that ConceptMaster significantly outperforms previous approaches, paving the way for generating personalized and semantically accurate videos across multiple concepts.
@@ 4 Comparison Methods Implementation
We supplement the implementation details of the compared methods for the MCVC task.
Comparing with naive solutions. We compare several open-sourced multi-concept image customization methods, including CustomDiffusion [30], SSR-Encoder [69], IP-Adapter [67] and MS-Diffusion [61], combining with the image-to-video (I2V) generation model I2VGen-XL [68], as a naive solution for the MCVC task with our ConceptMaster. While SSR-Encoder, IP-Adapter and MS-Diffusion do not need additional training and can be viewed as feed-forward customization models, CustomDiffusion requires users to manually few-shot examples for additional parameter training. Therefore, we collect the few-shot training samples from videos, where we randomly sample 1$\sim$5 frames as training images for each concept. Additionally, we follow its usage requirements, where we label specific character for each concept (i.e., s1* an elephant is playing with s2* a dog on the road). In addition to Tab. 1 and Fig. 5 in the main paper, we demonstrate more quantitative results in Fig. 10. Comparing with aforementioned methods, our end-to-end video customization models ConceptMaster could achieve better generation results more practicality than two-stage solutions.
Comparing with tuning-based methods. We compare the classical tuning-based method DreamBooth [50], where we segment the original videos into multiple clips for few-shot training examples and we train DreamBooth on our text-to-video models. We also follow its requirements to label specific character for each concept (i.e., s1* a man and s2* a woman dancing on a city street). In addition to Tab. 1 and Fig. 6 in the main paper, we demonstrate more quantitative results in Fig. 11. Comparing with DreamBooth, our ConceptMaster not only achieves significant advantages on both generated results and practicality, but also has better dynamic degree since DreamBooth also demonstrates poor dynamics under multi-concept scenarios.
@@ 5 More Discussions on Multi-Concept Embeddings Injection
We demonstrate more quantitative results between these three different multi-concept embeddings injection methods in Fig. 12. Our key insight is to inject the represented multi-concept embeddings into the diffusion models in a standalone cross-attention layer. While previous methods that the most representative ones include 1) BLIP-Diffusion [33], which combines visual and textual caption embeddings as the whole condition representation. 2) IP-Adapter [67], which encodes the whole image as visual embeddings and aggregates into models by a decoupled cross-attention layer. These two technical thoughts are widely adopted in previous image and video concept customization, which are different from our core insight. In addition to Tab. 1 and Fig. 5 in the main paper, we demonstrate more quantitative results in Fig. 12, where both these two methods can hardly deal with the identity decoupling problem with there are similar concepts (e.g., a man and a girl). Even when concepts have huge semantic differences (i.e., a woman and a dog), both two methods cannot maintain the concept fidelity of each concept. Additionally, when merging the visual and textual embeddings together as the conditions, the instruction following ability becomes very bad and the generated videos have the copy-paste effects. Therefore, our ConceptMaster adopts the most suitable manner of the injection of the multi-concept embeddings, which could represent and decouple multiple identities well.
@@ 6 More Discussions on Ablation Study
We demonstrate more quantitative results of the effectiveness of the Q-Former and DAM modules in Fig. 13. Initially, our ConceptMaster proposes to firstly utilizes a Q-Former network to integrate the dense visual tokens extracted by CLIP image encoder into the comprehensive visual embeddings. Additionally, we introduce the DAM module, which conducts the intra-pair attention module based on paired visual embeddings and textual descriptions representation. These designs demonstrate effective capability of decoupling and representing of multiple given references. While we also conduct several ablation architectures: 1) Replacing the Q-Former by an MLP layer, in Fig. 13 we can see that the visual appearances of provided man and red jacket could no longer be captured by the generation results. Therefore, the Q-Former is significant to assist the representation of comprehensive visual embeddings. 2) Removing the DAM module, where only the extracted visual embeddings are further injected into the diffusion models, and the textual descriptions are unused. Similar phenomenon could be observed in Fig. 13 as the concepts could not keep their fidelity. This is because the absence of adequate text label representation fails to effectively represent and differentiate the uniqueness of multiple concepts, and it also results in poor alignment with the original diffusion space. 3) Replacing DAM by firstly concatenating the visual and textual embeddings along channel dimension, and downsampling the dimension to the original one by an MLP layer. While the insufficient representation would lead to the inharmonious combination of multiple concepts. Since when concatenating along the channel dimension and then downsamling by MLP, the tokens could not conduct interaction among them. As in Fig. 13, the way that the man dressed in red jacket is unreasonable. 4) Replacing the intra-pair cross-attention by self-attention, where we firstly add the visual and textual embeddings and then conduct the self-attention operation. According to the qualitative and quantitative results in Tab. 3 and Fig. 8 in main paper and Fig. 13, the the cross-attention operation is better than self-attention, as the latter easily leads to more artifacts and inharmonious movements. Therefore, our proposed Q-Former and DAM modules would be the best designated architectures to simultaneously represent and decouple multiple references, and could create high-quality customized videos.
@@ 7 More Qualitative Results Demonstration
Figure 11: More Qualitative comparison on multi-concept customization between ConceptMaster and DreamBooth.

### v2

@@ 5.4 Multi-Concept Embeddings Injection Manner
In section 4.2, our ConceptMaster introduces a standalone MC-Injector to integrate the multi-concept visual-text aligned representation into the diffusion models. Some previous methods, represented as BLIP-Diffusion [34] and IP-Adapter [67], the former combines visual embeddings with textual caption embeddings as the whole condition representation, while the latter encodes the whole image as visual embeddings and aggregates into models by a decoupled cross-attention layer. However, in multi-concept scenarios, merging multi-modal features as the whole conditions can make it challenging to distinguish the semantic meanings among different identities. In addition, integrating all visual concepts on one image is not the optimal choice for multi-concept representation, especially when they contain similar visual appearances. We conduct the above two integration approaches of the multi-concept embeddings on our text-to-video generation models with the same training data for ConceptMaster. In Tab. 2 and Fig. 6, we can see that when customizing multiple concepts, both these two methods can hardly maintain the concept fidelity and deal with the identity decoupling problem. Additionally, when merging the textual and visual embeddings, the dynamic degree is significantly reduced, as the original text cross-attention layer is influenced by the additional visual embeddings. Therefore, our designed ConceptMaster adopts the optimal solution to inject the decoupled multi-concept embeddings into the diffusion models.
@@ 5.5 Ablation Study
In section 4.2, our ConceptMaster proposes to firstly utilizes a Q-Former network to integrate the dense visual tokens extracted by CLIP image encoder into the comprehensive visual embeddings. While after simply replacing the Q-Former by an MLP layer, the generated videos cannot capture the appearances of given images, as in Fig. 7, and the quantitative metrics also largely drop in Tab. 3.
In addition, in order to demonstrate the effectiveness of the designed DAM module, which is the intra-pair attention module based on paired visual embeddings and textual descriptions representation, we conduct several variants include: 1) Replacing the Q-Former, where we only use an MLP layer instead. 2) Removing the DAM module, where only the extracted visual embeddings are further injected into the diffusion models, and the textual descriptions are unused. 3) Replacing the intra-pair cross-attention by firstly concatenating the visual and textual embeddings along channel dimension (double the channel dimension), and integrating to the original channel dimension by an MLP layer. 4) Conducting the intra-pair self-attention instead of cross-attention, the features conduct self-attention are obtained by directly adding the visual and textual embeddings. In Tab. 3 and Fig. 7, we can see that the proposed DAM module is the optimal design. Cooperating the textual descriptions is significant, which not only enhances the uniqueness of each concept representation, but also assists the alignment of the multi-concept embeddings and the original diffusion model space. Furthermore, fusing visual and textual embeddings by the MLP layer is less comprehensive as it does not involve sufficient token-level interaction between them, lowering the concept representation and instruction-following in generated results. Additionally, the cross-attention operation is better than self-attention, which maintains consistent visual appearances in videos, while more artifacts would be created by the self-attention operation.
@@ 6 Conclusion
In this paper, we introduce ConceptMaster, an innovative framework that effectively addresses the critical issues of identity decoupling while maintaining concept fidelity when customizing multiple identities in videos. ConceptMaster introduces a novel strategy for learning decoupled multi-concept embeddings and injecting them into diffusion models in a standalone manner. This strategy ensures the quality of customized videos with multiple identities, even for highly similar visual concepts. To further address the scarcity of high-quality multi-concept video-entity data, we have established a meticulous data construction pipeline. This pipeline enables the systematic collection of precise multi-concept video-entity data across diverse concepts. Additionally, we have designed a comprehensive testing set to validate the effectiveness of our model from three critical dimensions across six different concept composition scenarios. Extensive experiments demonstrate that ConceptMaster significantly outperforms previous approaches, paving the way for generating personalized and semantically accurate videos across multiple concepts.
@@ 3 Discussions on Comparison between Our Data Collection Pipeline and Grounded-SAM
4 More details of Multi-Concept Evaluation Set
As mentioned in section 4.3, we manually collect reference images and give out suitable captions for these scenarios, in order to eliminate information leakage when extracting concepts from videos. We demonstrate the video caption templates for the six different scenarios in Tab. 4. The <ID> will be replaced with concept label.
@@ 5 Comparison Methods Implementation
We supplement the implementation details of the compared methods for the MCVC task. We compare several open-sourced multi-concept image customization methods, including CustomDiffusion [31], SSR-Encoder [69], IP-Adapter [67] and MS-Diffusion [61], combining with the image-to-video (I2V) generation model I2VGen-XL [68], as a naive solution for the MCVC task with our ConceptMaster. While SSR-Encoder, IP-Adapter and MS-Diffusion do not need additional training and can be viewed as feed-forward customization models, CustomDiffusion requires users to manually few-shot examples for additional parameter training. Therefore, we use the collected reference images for each concept as training samples. Additionally, we follow its usage requirements, where we label specific character for each concept (i.e., s1* elephant is playing with s2* dog on the road). In addition to Tab. 1 and Fig. 5 in the main paper, we demonstrate more quantitative results in Fig. 9. Comparing with aforementioned methods, our end-to-end video customization models ConceptMaster could achieve better generation results more practicality than two-stage solutions.
@@ 6 More Discussions on Multi-Concept Embeddings Injection
We demonstrate more quantitative results between these three different multi-concept embeddings injection methods in Fig. 10. Our key insight is to inject the represented multi-concept embeddings into the diffusion models in a standalone cross-attention layer. While previous methods that the most representative ones include 1) BLIP-Diffusion [34], which combines visual and textual caption embeddings as the whole condition representation. 2) IP-Adapter [67], which encodes the whole image as visual embeddings and aggregates into models by a decoupled cross-attention layer. These two technical thoughts are widely adopted in previous image and video concept customization, which are different from our core insight. In addition to Tab. 1 and Fig. 5 in the main paper, we demonstrate more quantitative results in Fig. 10, where both these two methods can hardly deal with the identity decoupling problem when there are similar concepts (e.g., a man and a girl). Even when concepts have huge semantic differences (i.e., a man and a red jacket), both two methods cannot maintain the concept fidelity of each concept. Additionally, when merging the visual and textual embeddings together as the conditions, the instruction following ability becomes unsatisfactory and the dynamic degree also drops. Therefore, our ConceptMaster adopts the most suitable manner of the injection of the multi-concept embeddings, which could represent and decouple multiple identities well.
@@ 7 More Discussions on Ablation Study
We demonstrate more quantitative results of the effectiveness of the Q-Former and DAM modules in Fig. 11. Initially, our ConceptMaster proposes to firstly utilizes a Q-Former network to integrate the dense visual tokens extracted by CLIP image encoder into the comprehensive visual embeddings. Additionally, we introduce the DAM module, which conducts the intra-pair attention module based on paired visual embeddings and textual descriptions representation. These designs demonstrate effective capability of decoupling and representing of multiple given references. While we also conduct several ablation architectures: 1) Replacing the Q-Former by an MLP layer, in Fig. 11 we can see that the visual appearances of provided woman and dog could no longer be captured by the generation results. Therefore, the Q-Former is significant to assist the representation of comprehensive visual embeddings. 2) Removing the DAM module, where only the extracted visual embeddings are further injected into the diffusion models, and the textual descriptions are unused. Similar phenomenon could be observed in Fig. 11 as the concepts could not keep their fidelity. This is because the absence of adequate text label representation fails to effectively represent and differentiate the uniqueness of multiple concepts, and it also results in poor alignment with the original diffusion space. 3) Replacing DAM by firstly concatenating the visual and textual embeddings along channel dimension, and downsampling the dimension to the original one by an MLP layer. While the insufficient representation would lead to the inharmonious combination of multiple concepts. Since when concatenating along the channel dimension and then downsamling by MLP, the tokens could not conduct interaction among them. As in Fig. 11, the way that the woman who walks dog on the beach is unreasonable. 4) Replacing the intra-pair cross-attention by self-attention, where we firstly add the visual and textual embeddings and then conduct the self-attention operation. According to the qualitative and quantitative results in Tab. 3 and Fig. 7 in main paper and Fig. 11, the cross-attention operation is better than self-attention, as the latter easily leads to more artifacts and inharmonious movements. Therefore, our proposed Q-Former and DAM modules would be the best designated architectures to simultaneously represent and decouple multiple references, and could create high-quality customized videos.
@@ 8 More Qualitative Results Demonstration
Diverse Scenarios Caption Templates
1) Multiple Persons <ID1> and <ID2> hugging each other in front of a bridge.
<ID1> and <ID2> kissing each other in front of a bridge.
<ID1> and <ID2> walking down a city street.
<ID1> and <ID2> dancing on a city street.
<ID1> and <ID2> smiling and shaking hands in the office.
<ID1> and <ID2> walking on the beach.
2) Persons with Livings <ID1> walking <ID2> on the beach.
<ID1> walking <ID2> in the woods.
<ID1> petting <ID2> in the park.
<ID1> feeding <ID2> in the garden.
<ID1> and <ID2> running on the grass.
<ID1> rides <ID2> running on the farm.
<ID1> petting <ID2> in the stable.
<ID1> raising <ID2> in the garden.
<ID1> holding <ID2> and walking on the street.
<ID1> feeding <ID2> on the street.
<ID1> and <ID2> walking in the desert.
3) Persons with Stuffs <ID1> wearing <ID2> walking in the shopping mall.
<ID1> wearing <ID2> walking along the river.
<ID1> wearing <ID2> running in the stadium.
<ID1> wearing <ID2> dancing on the floor.
<ID1> rides <ID2> in the desert.
<ID1> rides <ID2> on the road.
<ID1> is happily playing <ID2> on the bench.
<ID1> is happily playing <ID2> in the desert.
<ID1> taking <ID2>, snowy winter.
<ID1> holding <ID2> and walking in the park.
<ID1> raising <ID2> in the garden.
<ID1> holding <ID2> and walking on the street.
<ID1> walking around <ID2> on the street.
<ID1> dancing in front of <ID2>, snowy day.
<ID1> walking in front of <ID2> on the street.
<ID1> dancing in front of <ID2> on the street.
4) Multiple Livings <ID1> is playing with <ID2> on the road.
<ID1> and <ID2> walking on the grass.
<ID1> walking around <ID2> in the desert.
<ID1> is playing with <ID2> on the street.
5) Livings with Stuffs <ID1> walking around <ID2> in a grassy field.
<ID1> playing with <ID2> in a grassy field.
<ID1> walking around <ID2> at home.
<ID1> is playing with <ID2> at home.
<ID1> is walking around <ID2> on the road.
<ID1> is staying besides <ID2> in the snow.
<ID1> is drinking water from <ID2> in the garden.
<ID1> is drinking water from <ID2> on the road.
6) Persons with both Livings and Stuffs <ID1> walking <ID2> in front of <ID3> on the street.
<ID1> and <ID2> walking around <ID3> in the snow.
<ID1> and <ID2> walking around <ID3> on the road.
<ID1> is happily playing <ID2> and <ID3> surrounds in the garden.
<ID1> taking <ID2> and walking with <ID3> on the grass.
Table 4: Caption templates for Multi-Concept Evaluation Set.